# Thunder-NUBench: A Benchmark for LLMs’ Sentence-Level Negation Understanding

### v3

@@ Appendix G HoVer Dataset
The statement to be verified, often requiring multi-article evidence
@@
List of Wikipedia article titles and sentence indices providing evidence
@@
HoVer is constructed on top of the HotpotQA dataset, which is designed to evaluate multi-hop reasoning in question answering. HotpotQA itself is a large-scale collection of Wikipedia-based QA pairs created to address the limitations of prior QA datasets, which often fail to require complex reasoning or explanatory answers (yang2018hotpotqa). The construction of HoVer involves rewriting HotpotQA question-answer pairs into claim statements, which are then validated and labeled by annotators. Claims are extended to require multi-hop evidence from up to four Wikipedia articles and are systematically modified to increase complexity. Final labels are assigned as SUPPORTED or NOT-SUPPORTED (jiang2020hover).
@@ Appendix H Wikipedia Summary Dataset
description A brief description or category for the article (when available). French mathematician
@@
The extracted summary or introduction section of the article,
typically more concise than the full text.
Alain Connes (; born 1 April 1947) is a French mathematician…
@@
The complete article text (when included), encompassing the
full body of the Wikipedia page.
Alain Connes (; born 1 April 1947) is a French mathematician…
@@
The Wikipedia Summary Dataset contains the titles and introductory summaries of English Wikipedia articles, extracted in September 2017. A summary or introduction in this context refers to the content from the article title up to the content outline (i.e., before the first section heading). The dataset was originally released via GitHub [^2] https://github.com/tscheepers/Wikipedia-Summary-Dataset, but is now accessible through the Hugging Face Hub [^3] https://huggingface.co/datasets/jordiclive/wikipedia-summary-dataset. The dataset license is not explicitly mentioned, but as the original Wikipedia data is distributed under the CC BY-SA 4.0, it is assumed that the dataset would be distributed under the same license. For licensing details, refer to the Wikimedia Terms of Use  [^4] https://foundation.wikimedia.org/wiki/Policy:Terms_of_Use. Table 9 offers an overview of the dataset’s structure. The dataset comprises approximately 430,000 articles, only providing the training set (scheepers2017compositionality).
@@ Appendix I Human Review Protocol
Overall, this iterative and layered procedure ensures that every instance in the multiple-choice dataset is independently reviewed multiple times, leading to stable guidelines and a consistent dataset.
@@
To construct the sentence-negation pair dataset, we begin by randomly sampling sentences labeled as "supported facts" from the HoVer dataset. Since the original data often contains grammatical errors, we utilize OpenAI’s API (OpenAI) to automatically correct these issues. In cases where the selected text consists of multiple sentences, we merge or split them as needed to ensure that each example is a single sentence, aligning with our sentence-level task objective.
@@ 8 Paraphrased sentence:
Appendix L Thunder-NUBench Dataset Structure
L.1 Sentence-Negation Pair Dataset
  • index: the index of the data.
  • premise: the original sentence.
  • hypothesis: its logically negated form.
L.2 Multiple-Choice Dataset
  • index: the index of the data.
  • sentence: the original sentence.
  • choice1: standard negation (correct answer).
  • choice2: local negation (subclausal negation).
  • choice2_type: specifies the type of local negation.
  • choice4: paraphrase (semantically equivalent).
@@
Appendix M Incorporating Thunder-NUBench into LM Evaluation Harness
This section describes how Thunder-NUBench is integrated into the LM Evaluation Harness (eval-harness) for zero-shot and few-shot evaluation, both in completion-based and option-selection settings.
1task: nubench_completion
2dataset_path: {dataset_path}/NUBench
3dataset_name: multiple-choice
4output_type: multiple_choice
5test_split: test
6fewshot_split: demonstration
7process_docs: !function utils.process_docs_completion
8doc_to_text: "{{query}}"
9doc_to_target: "{{gold}}"
10doc_to_choice: "choices"
11metric_list:
12 - metric: acc
13 aggregation: mean
14 higher_is_better: true
Listing 7: NUBench/NUBench_completion.yaml
1task: nubench_option
2dataset_path: {dataset_path}/NUBench
3dataset_name: multiple-choice
4output_type: generate_until
5test_split: test
6fewshot_split: demonstration
7process_docs: !function utils.process_docs_option
8doc_to_text: "{{query}}"
9doc_to_target: "{{answerLetter}}"
10generation_kwargs:
11 until:
12 - "</s>"
[…]
14metric_list:
15 - metric: exact_match
16 aggregation: mean
17 higher_is_better: true
18 ignore_punctuation: true
19 ignore_case: true
20filter_list:
21 - name: get_response
22 filter:
23 - function: "regex"
[…]
25 - function: remove_whitespace
26 - function: "regex"
[…]
28 - function: take_first
29dataset_kwargs:
30 trust_remote_code: true
Listing 8: NUBench/NUBench_option.yaml
1import re
2import datasets
3import random
4
5def process_docs_completion(dataset: datasets.Dataset) -> datasets.Dataset:
6 def _process_doc(doc):
[…]
8
9 if doc.get("choice2_type", "") == "non-applicable":
10 choices = [doc["choice1"], doc["choice3"], doc["choice4"]]
11 else:
12 choices = [doc["choice1"], doc["choice2"], doc["choice3"], doc["choice4"]]
13
14 return {
15 "query": prompt,
16 "choices": choices,
17 "gold": 0
18 }
19 return dataset.map(_process_doc)
20
21
22def process_docs_option(dataset: datasets.Dataset, seed: int = 42) -> datasets.Dataset:
23 rng = random.Random(seed)
24
25 def _process_doc(doc):
[…]
27
28 c1 = doc.get("choice1", "")
29 c2 = doc.get("choice2", "")
30 c3 = doc.get("choice3", "")
31 c4 = doc.get("choice4", "")
32 c2_na = str(doc.get("choice2_type", "")).lower() == "non-applicable"
33
34 gathered = []
35 keys = []
36
37 if c1:
38 gathered.append(c1); keys.append("choice1")
39 if c2 and not c2_na:
40 gathered.append(c2); keys.append("choice2")
41 if c3:
42 gathered.append(c3); keys.append("choice3")
43 if c4:
44 gathered.append(c4); keys.append("choice4")
45
46 assert len(gathered) >= 3, "Need at least 3 choices for multiple-choice."
47
48 paired = list(zip(gathered, keys))
49 rng.shuffle(paired)
50 choices, choice_keys = zip(*paired)
51
52 try:
53 answer_idx = choice_keys.index("choice1")
54 except ValueError:
55 answer_idx = 0
56
57 label_map = ["A", "B", "C", "D"]
58 labels = label_map[:len(choices)]
59
60 lines = [
61 "Given the following instruction and candidate answers, choose the single best answer.",
62 f"Instruction: {initial_prompt}",
63 ]
64 for lab, ch in zip(labels, choices):
65 lines.append(f"{lab}. {ch}")
66 joined = ", ".join(labels)
67 lines += ["", f"Your response should be one of {joined}.", "Only output the letter.", "Answer:"]
[…]
69
70 return {
71 "query": prompt,
72 "choices": list(choices),
73 "choice_keys": list(choice_keys),
74 "answerKey": answer_idx,
75 "answerLetter": labels[answer_idx],
76 }
77
78 return dataset.map(_process_doc)
Listing 9: NUBench/utils.py
Appendix N Finetuning via LLaMA-Factory
We detail our supervised fine-tuning setup using LLaMA-Factory (zheng2024llamafactory) with LoRA (hu2022lora) on Thunder-NUBench training data, including configuration of the fine-tuning and instruction-based examples in Alpaca format (alpaca).
The YAML configuration provided in Listing 10 is specific to the LLaMA-3.1-8B model. Other models (e.g., Qwen or Mistral) can be fine-tuned similarly by modifying the model_name_or_path and template fields in the configuration file accordingly.
1### model
2model_name_or_path: Llama-3.1-8B
3trust_remote_code: true
4
5### method
6stage: sft
7do_train: true
8finetuning_type: lora
9lora_rank: 8
10lora_target: all
11
12### dataset
13dataset: nubench_train
14template: llama3
15cutoff_len: 512
16max_samples: 5000
17overwrite_cache: true
18preprocessing_num_workers: 16
19dataloader_num_workers: 4
20
21### output
22output_dir: lora/sft/Llama-3.1-8B
23logging_steps: 10
24save_steps: 500
25plot_loss: true
26overwrite_output_dir: true
27save_only_model: false
28report_to: none # choices: [none, wandb, tensorboard, swanlab, mlflow]
29
30### train
31per_device_train_batch_size: 1
32gradient_accumulation_steps: 8
33learning_rate: 1.0e-4
34num_train_epochs: 3.0
35lr_scheduler_type: cosine
36warmup_ratio: 0.1
37bf16: true
38ddp_timeout: 180000000
39resume_from_checkpoint: null
Listing 10: Llama-3.1-8B_lora_sft.yaml
1[
2 {
3 "instruction": "Negate the sentence.",
4 "input": "Sentence: Eddie Vedder was born before Nam Woo-hyun.",
5 "output": "Eddie Vedder wasnt born before Nam Woo-hyun."
6 },
7 {
8 "instruction": "Negate the sentence.",
9 "input": "Sentence: Halestorm is from Pennsylvania, while Say Anything is from California.",
10 "output": "Halestorm is not from Pennsylvania, while Say Anything is from California."
11 },
12 (...)
13]
Listing 11: Sentence-Negation Pair dataset for training in alpaca format.
Appendix O Prompt Selection for In-Context Learning
We explore a range of prompt types for in-context learning, from minimal instructions to more detailed variants (zhao2021calibrate; li2023practical; wan2023better). Specifically, we evaluate three prompt styles:
@@
    Simple prompt: a minimal instruction, “Negate the sentence.”
@@
    Definition prompt: a concise description of the task, “Negate the main predicate of the main clause so that the proposition is logically reversed.”
  3. 3.
    Detail prompt: an extended instruction with explicit step-by-step guidelines on identifying the main predicate, preserving other sentence elements, handling antonyms, and applying negation consistently across logical operators (see Listing 12 for the full format).
"Reverse the truth value of the main predicate (verbal phrase) in the main clause, while preserving all other elements of the main clause unchanged.
@@
4) If the sentence contains multiple propositions connected by logical operators (e.g., and, or, conditional constructions), negate it in a way that reverses the entire proposition (e.g., A and B => not A or not B; If A then B => A and not B).
Sentence: {doc[’sentence’]}"
Listing 12: Detail prompt format.
Across both completion-based and option-selection evaluation settings, we found that the simple prompt consistently achieved the highest average performance in zero-shot and few-shot settings (See results in Appendix P.) While the definition-based and detail prompts occasionally provided more explicit guidance, they did not improve performance overall. Therefore, we adopt the simple prompt as our default setting in the main experiments. The complete prompt formats, including prompt–response structures for each evaluation setting, are provided in Listings 13 and 14, based on the simple prompt.
Prompt] "Negate the sentence.
@@
Listing 13: Completion-based Format.
@@
Instruction: Negate the sentence.
@@
C. Chromosome 2 is the smallest human chromosome, spanning fewer than 50 million base pairs and representing less than two percent of the total DNA in human cells.\
@@
Listing 14: Option-selection Format.
@@
Appendix P Total Model Performance on Thunder-NUBench
@@
evaluation
setting
comple-
tion
option
comple-
tion
@@
comple-
tion
option
gemma-2b 0.437 0.197
0.437
@@
0.540
(±0.011)
0.557
(±0.006)
0.526
(±0.006)
0.587
@@
0.532
(±0.005)
0.745 0.604
Llama-3.1-8B 0.439 0.459
0.493
(±0.007)
0.549
(±0.012)
0.596
(±0.006)
0.635
(±0.008)
0.635
(±0.006)
0.667
(±0.011)
0.797 0.556
Llama-3.1-8B-
Instruct
0.512 0.744
0.571
(±0.007)
0.708
(±0.007)
0.646
(±0.007)
0.726
(±0.008)
0.675
(±0.004)
0.729
(±0.011)
0.756 0.747
Mistral-7B-v0.3 0.425 0.376
0.471
(±0.003)
0.603
(±0.008)
0.548
(±0.012)
0.702
(±0.010)
0.597
(±0.002)
0.720
(±0.013)
0.773 0.386
Mistral-7B-
Instruct-v0.3
0.619 0.642
0.635
(±0.007)
@@
0.761 0.656
Qwen2.5-7B-
Instruct
0.542 0.676
0.558
(±0.007)
0.637
(±0.005)
0.620
(±0.004)
0.623
(±0.005)
0.659
(±0.005)
0.614
(±0.005)
0.718 0.793
gpt-4o-mini - 0.643 -
0.682
(±0.002)
0.687
(±0.005)
0.704
(±0.008)
gpt-4.1-mini - 0.797 -
0.822
(±0.005)
0.841
(±0.008)
0.846
(±0.005)
claude-3-5-
haiku-latest
- 0.722 -
0.505
(±0.016)
0.476
(±0.009)
0.633
(±0.004)
2-3B models
average
0.456 0.404 0.484 0.446 0.556 0.464 0.585 0.469 0.731 0.428
7-8B models
average
0.517 0.569 0.545 0.631 0.609 0.672 0.649 0.682 0.769 0.655
api models
average
- 0.721 - 0.670 - 0.668 - 0.728 - -
base models
average
0.448 0.408 0.470 0.513 0.547 0.564 0.591 0.582 0.760 0.507
instruct-tuned
models average
0.533 0.589 0.567 0.59 0.625 0.601 0.652 0.599 0.746 0.608
total average 0.491 0.538 0.518 0.573 0.586 0.598 0.622 0.615 0.753 0.558
Table 11: Zero-shot, few-shot, and SFT evaluation results on Thunder-NUBench with the simple prompt. SD denotes standard deviation across random seeds or runs. Few-shot results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and one, five, and ten demonstration examples (1-, 5-, 10-shot). Red text indicates the model with the highest performance in each column (excluding API models).
@@
evaluation
setting
comple-
tion
option
comple-
tion
@@
comple-
tion
option
gemma-2b 0.443 0.176
0.427

### v4

@@ Appendix G HoVer Dataset
The statement to be verified,
often requiring multi-article evidence
@@
List of Wikipedia article titles and sentence indices
providing evidence
@@
HoVer is constructed on top of the HotpotQA dataset, which is designed to evaluate multi-hop reasoning in question answering. HotpotQA itself is a large-scale collection of Wikipedia-based QA pairs created to address the limitations of prior QA datasets, which often fail to require complex reasoning or explanatory answers (Yang et al., 2018). The construction of HoVer involves rewriting HotpotQA question-answer pairs into claim statements, which are then validated and labeled by annotators. Claims are extended to require multi-hop evidence from up to four Wikipedia articles and are systematically modified to increase complexity. Final labels are assigned as SUPPORTED or NOT-SUPPORTED (Jiang et al., 2020).
@@ Appendix H Wikipedia Summary Dataset
description
A brief description or category for the article
(when available).
French mathematician
@@
The extracted summary or introduction section
of the article, typically more concise than the full text.
Alain Connes (; born 1 April 1947)
is a French mathematician…
@@
The complete article text (when included),
encompassing the full body of the Wikipedia page.
Alain Connes (; born 1 April 1947)
is a French mathematician…
@@
The Wikipedia Summary Dataset contains the titles and introductory summaries of English Wikipedia articles, extracted in September 2017. A summary or introduction in this context refers to the content from the article title up to the content outline (i.e., before the first section heading). The dataset was originally released via GitHub [^3] https://github.com/tscheepers/Wikipedia-Summary-Dataset, but is now accessible through the Hugging Face Hub [^4] https://huggingface.co/datasets/jordiclive/wikipedia-summary-dataset. The dataset license is not explicitly mentioned, but as the original Wikipedia data is distributed under the CC BY-SA 4.0, it is assumed that the dataset would be distributed under the same license. For licensing details, refer to the Wikimedia Terms of Use  [^5] https://foundation.wikimedia.org/wiki/Policy:Terms_of_Use. Table 9 offers an overview of the dataset’s structure. The dataset comprises approximately 430,000 articles, only providing the training set (Scheepers, 2017).
@@ Appendix I Human Review Protocol
Overall, this iterative and layered procedure ensures that every instance in the multiple-choice dataset is independently reviewed across multiple stages, leading to stable guidelines and a consistent dataset.
Edit-based reliability and quality control statistics.
Because our review process is revision-based rather than label-based, standard inter-annotator agreement metrics (e.g., Cohen’s $\kappa$) are not directly applicable. Instead of assigning categorical labels, reviewers have directly edited or deleted problematic instances to enforce the dataset guidelines.
We therefore report edit-based quality control statistics as a proxy for annotation reliability, including (i) the number of instances escalated to adjudication meetings and (ii) the number of instances that required edits after cross-checking. All cases requiring edits have been discussed in meetings and resolved by consensus, ensuring consistency across the dataset.
Subset # items
# Brought
to meeting
# Edited
Edit
rate
Sentence-
Negation
Standard
Negation
3,992 469 72 1.80%
Multiple Choice
Standard/
Local Negation
1,102 126 34 3.09%
Contradiction/
Paraphrase
1,102 409 117 10.62%
Implication
(add-on)
201 49 16 7.96%
Table 10: Edits include rewriting or deleting instances that violate guidelines.
Overall, the relatively low edit rates after cross-checking indicate that most instances satisfied the guidelines upon independent review, while meeting-based adjudication played a critical role in resolving ambiguous or complex cases.
@@
To construct the sentence-negation pair dataset, we begin by randomly sampling sentences labeled as "supported facts" from the HoVer dataset. Since the original data often contains grammatical errors, we utilize OpenAI’s API (OpenAI, 2025) to automatically correct these issues. In cases where the selected text consists of multiple sentences, we merge or split them as needed to ensure that each example is a single sentence, aligning with our sentence-level task objective.
@@ 8 Paraphrased sentence:
Appendix L Thunder-NUBench Dataset Structure
L.1 Sentence-Negation Pair Dataset
  • index: the index of the data.
  • premise: the original sentence.
  • hypothesis: its logically negated form.
L.2 Multiple-Choice Dataset
  • index: the index of the data.
  • sentence: the original sentence.
  • choice1: standard negation (correct answer).
  • choice2: local negation (subclausal negation).
  • choice2_type: specifies the type of local negation.
  • choice4: paraphrase (semantically equivalent).
@@
Appendix M Prompt Templates Used in Evaluation
We design and employ two instruction formats for in-context learning. These instructions differ in the level of guidance they provide, ranging from a task definition to a detailed procedural instruction. Specifically, we use the following two instruction styles. These instruction formats correspond to the task instruction component shown in Figure 1.
@@
    Definition instruction: A concise definition of standard negation, which specifies the target operation,
    "Standard negation is sentential negation that reverses the truth value of the sentence by negating the main predicate(s) of the main clause(s). Keep the rest of the sentence content unchanged."
@@
    Detailed instruction: A step-by-step instruction that presents standard negation as an explicit procedure. This prompt describes how to identify the main clause and its main predicate, preserve other sentence elements, apply syntactic negation or complementary antonyms where appropriate, and negate complex sentences by reversing their logical structure. The full prompt format is shown in Listing 7.
Standard negation reverses the truth value of the main predicate in the main clause while keeping all other elements of the main clause unchanged.
Do not negate subordinate clauses or modify other parts of the sentence.
To do this:
@@
4) If the sentence contains multiple propositions connected by logical operators (e.g., and, or, conditional constructions), negate it in a way that reverses the entire proposition (e.g., A and B -> not A or not B; If A then B -> A and not B).
Listing 7: Detailed instruction format.
Below, we show the prompt formats used for completion-based and option-selection evaluations, including prompt–response structures for each evaluation setting.
Prompt] "{instruction format}
Generate the standard negation of the given sentence.
@@
Listing 8: Completion-based Format.
@@
{instruction format}
Generate the standard negation of the given sentence.
@@
C. Chromosome 2 is the smallest human chromosome, spanning fewer than 50 million base pairs and representing less than two percent of the total DNA in human cells.
@@
Listing 9: Option-selection Format.
Appendix N Evaluation Setup
We use 8 NVIDIA GeForce RTX 3090 GPUs with 24GB of memory and CUDA 12.4 for all fine-tuning and evaluation.
N.1 Models
We evaluate a diverse set of pretrained and instruction-tuned models across multiple model families and scales, ranging from about a billion to 14 billion parameters. Our experiments include models from Gemma (Team et al., 2024), Qwen (Yang et al., 2025), LLaMA (Grattafiori et al., 2024), and Mistral families (Jiang et al., 2023), as well as API-based models from OpenAI (Achiam et al., 2023; Hurst et al., 2024) and Anthropic (Anthropic, 2024). API models are evaluated only in the option-selection setting, as they do not provide access to token-level log probabilities required for completion-based evaluation.
  • Gemma models:
    • gemma-2-2b, gemma-2-2b-it
    • gemma-2-9b, gemma-2-9b-it
  • Qwen models:
    • Qwen3-0.6B-Base, Qwen3-0.6B
    • Qwen3-1.7B-Base, Qwen3-1.7B
    • Qwen3-4B-Base, Qwen3-4B
    • Qwen3-8B-Base, Qwen3-8B
    • Qwen3-14B-Base, Qwen3-14B
    • Qwen3-32B
  • LLaMA models:
    • Llama-3.1-8B, Llama-3.1-8B-Instruct
    • Llama-3.2-1B, Llama-3.2-1B-Instruct
    • Llama-3.2-3B, Llama-3.2-3B-Instruct
  • Mistral models:
    • Mistral-7B-v0.3, Mistral-7B-Instruct-v0.3
    • Mistral-Nemo-Base-2407 (12B size), Mistral-Nemo-Instruct-2407 (12B size)
    • Mistral-Small-24B-Base-2501, Mistral-Small-24B-Instruct-2501
  • API models:
    • GPT: GPT-4o mini, GPT-4o, GPT-4.1 mini, GPT-4.1
    • Claude: Haiku 4.5, Sonnet 4.5
N.2 Zero-shot and few-shot evaluation
For each model, we evaluate performance in both zero-shot and few-shot settings using the Language Model Evaluation Harness (Gao et al., 2024). In the few-shot scenario, we use examples from the demonstration set as in-context demonstrations. Results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and are reported for one, five, and ten examples from the demonstration set (1-shot, 5-shot, and 10-shot). We present the performance results on the test set for each model and prompt configuration.
N.3 Supervised fine-tuning
We conduct Supervised Fine-Tuning (SFT) using the LLaMA-Factory framework (Zheng et al., 2024) on the Sentence-Negation Pair dataset from Thunder-NUBench. The dataset is formatted in the Alpaca instruction style (Taori et al., 2023). To achieve parameter-efficient training, we apply Low-Rank Adaptation (LoRA) (Hu et al., 2022) with a rank of 8, targeting all linear layers. The fine-tuning process is carried out for three epochs, using a batch size of 1, a gradient accumulation step of 8, cosine learning rate scheduling, and bfloat16 precision. We apply supervised fine-tuning only to models with fewer than 10B parameters, in order to focus on analyzing the effects of SFT on relatively smaller models, where instruction tuning is expected to have a more pronounced impact. After the SFT, we evaluate the model’s zero-shot performance to directly assess its ability to generalize from instruction tuning without being influenced by in-context examples. It is important to note that API models are not fine-tuned, as they are not compatible with the LLaMA-Factory framework.
Appendix O Total Model Performance on Thunder-NUBench
This section provides the full set of results on Thunder-NUBench for all evaluated models and training settings. Following the prompt templates in Appendix M, we report results separately for the definition instruction and detailed instruction styles.
For each instruction style, models are evaluated under three conditions: (1) zero-shot, (2) few-shot with 1, 5, and 10 in-context demonstrations (averaged across random seeds), and (3) supervised fine-tuning (SFT). For SFT, models are evaluated in the zero-shot setting to isolate the effect of task-specific training without the influence of in-context examples.
Tables 14 and 15 report the results of open-weight models under the definition and detailed instruction styles, respectively. Tables 16 and 17 present the corresponding results for API models under the same instruction settings.
Appendix P General Benchmark Performance after SFT
To assess whether supervised fine-tuning (SFT) on Thunder-NUBench affects performance on broader natural language understanding tasks, we evaluate models on six widely used benchmarks: ARC-Challenge, ARC-Easy, GSM8K, HellaSwag, OpenBookQA, and WinoGrande (Clark et al., 2018; Cobbe et al., 2021; Zellers et al., 2019; Mihaylov et al., 2018; Sakaguchi et al., 2021). These datasets cover a diverse range of domains, including commonsense reasoning, scientific knowledge, mathematics, and reading comprehension.
Table 18 and Table 19 summarize the results. We find that performance on general benchmarks remains broadly stable after SFT, indicating that fine-tuning on Thunder-NUBench does not substantially harm general capabilities.
@@
Appendix Q Total Error Analysis of Model Predictions on the Thunder-NUBench
This section extends the error analysis presented in Section 5.3, providing the complete results. In particular, we examine (i) incorrect choice distributions and (ii) confusion rates for local negation categories, comparing models of different sizes and training paradigms (pretrained vs. instruction-tuned) under both zero-shot and few-shot conditions.
definition detailed
completion option completion option
Gemma2 Table 20 Table 21 Table 22 Table 23
Qwen3 Table 24 Table 25 Table 26 Table 27
Llama3 Table 28 Table 29 Table 30 Table 31
Mistral Table 32 Table 33 Table 34 Table 35
API Table 36 Table 37
Table 12: Complete prediction analysis by model family.
We report few-shot results using a fixed random seed (1234). Averaging over multiple seeds was avoided, as it could obscure specific error patterns and make confusion analysis less interpretable. Each table follows the same instruction format, reporting error rates, incorrect choice distributions, and confusion rates across local negation subcategories under zero-shot, few-shot, and SFT conditions.
For API models, we also track cases labeled as “Answer Format Wrong.” This category captures instances where the model’s output does not follow the required answer format (selecting strictly one of A, B, C, or D). Because such responses cannot be mapped to a specific incorrect option, they are not included in the incorrect choice distribution or confusion rate. Instead, we report their raw counts alongside the other error statistics. This also serves as an indicator of the model’s ability to follow output-format instructions.
Appendix R Human Evaluation
The human evaluation was approved by the Institutional Review Board (IRB No. 2512/004-015). All participants provided informed consent through an official IRB-approved consent form prior to participation and were fully informed of the study’s purpose, procedures, potential risks and benefits, and their right to withdraw at any time without penalty. Participants were also notified that only aggregate statistics of their evaluation scores would be reported.
All responses were collected anonymously, and no personally identifiable information was included in the dataset. Participants were compensated KRW 50,000 for a two-hour session, which exceeds the minimum hourly wage (KRW 10,320 per hour) and constitutes adequate compensation. To minimize fatigue and ensure reliable responses, participants were given sufficient rest during the evaluation, and the total duration was limited to approximately two hours per session. Participants were required to be native or highly proficient English users (CEFR [^7] Common European Framework of Reference for languages. C-level or above), as verified during recruitment. A total of 11 adult participants aged in their 20s to 30s took part in the study, including 6 highly proficient English users at the CEFR C level and 5 native or dominant English speakers.
A total of 50 items were randomly sampled from the multiple-choice evaluation split of Thunder-NUBench, with stratified sampling to preserve the distribution of distractor types. Specifically, the sample included 12 items each for relative clause, participle clause, compound sentence, and adverbial clause (local negation type), and 2 items without local negation. All items were presented using the detailed instruction format. Participants were asked to select the option corresponding to standard negation for each item. No example items were provided during the evaluation.
Mean Median Min Max SD
0.78 0.82 0.58 0.96 0.139
Table 13: Human evaluation results.
Table 13 summarizes the results of the human evaluation conducted with 11 participants. Human accuracy ranges from 0.58 to 0.96, with a mean of 0.78. This range indicates that distinguishing standard negation from closely related alternatives can be challenging even for human readers, consistent with prior observations that negation processing requires careful semantic and syntactic reasoning. The average human accuracy is higher than the zero-shot performance of most models reported in this paper, suggesting that humans generally handle sentence-level negation more reliably without task-specific training.
@@
evaluation setting completion option
completion
@@
completion option
gemma-2-2b 0.436 0.236
0.480
(±0.006)
0.328
(±0.004)
0.565
(±0.010)
0.373
(±0.009)
0.599
(±0.004)
0.398
(±0.011)
0.781 0.260
gemma-2-2b-it 0.516 0.504
0.514
(±0.008)
0.501
(±0.008)
0.569
(±0.007)
0.510
(±0.012)
0.594
(±0.003)
0.540
(±0.007)
0.741 0.543
gemma-2-9b 0.470 0.468
0.499
@@
0.715
(±0.008)
0.650
(±0.008)
0.793 0.692
Mistral-Nemo-
Base-2407 (12B)
0.460 0.412
0.490
(±0.010)
0.520
(±0.007)
0.600
(±0.003)
0.568
(±0.008)
0.649
(±0.007)
0.599
(±0.009)
Mistral-Nemo-
Instruct-2407 (12B)
0.487 0.604
0.528
@@
total average 0.464 0.519 0.499 0.559 0.577 0.599 0.615 0.622 0.740 0.554
Table 14: Zero-shot, few-shot, and SFT evaluation results on Thunder-NUBench with the definition instruction. SD denotes standard deviation across random seeds or runs. Few-shot results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and one, five, and ten demonstration examples (1-, 5-, 10-shot). Red text indicates the model with the highest performance in each column.
@@
evaluation setting completion option
completion
@@
completion option
gemma-2-2b 0.438 0.276
0.494
(±0.005)
0.343
(±0.005)
0.570
(±0.004)
0.360
(±0.018)
0.604
(±0.004)
0.378
(±0.010)
0.768 0.259
gemma-2-2b-it 0.505 0.470
0.488
(±0.013)
0.474
(±0.009)
0.545
(±0.007)
0.487
(±0.011)
0.580